For two numbers whose sum is q and whose difference is r, write an expression for each number in terms of q and r.
Answer by sdmmadam@yahoo.com(530)

For two numbers whose sum is q and whose difference is r, write an expression for each number in terms of q and r.
Let the numbers be M and N say with M >N
sum is q and difference is r
That is M+N = q ----(1)
and M-N= r ----(2)
(1)+(2) implies
(M+M) = (q+r)
2M = (q+r)
Therefore M= (q+r)/2 ----(*)
Putting (*) in (1)
N = q-M
=q-(q+r)/2
=[2q-(q+r)]/2
=(2q-q-r)/2
=(q-r)/2
Answer: The numbers are (q+r)/2 and (q-r)/2
Verification: Obviously their addition gives q and difference gives r
Hence our values are correct